# Towards Robust Model-Based Reinforcement Learning Against Adversarial Corruption

### v2

@@ Abstract
This study tackles the challenges of adversarial corruption in model-based reinforcement learning (RL), where the transition dynamics can be corrupted by an adversary. Existing studies on corruption-robust RL mostly focus on the setting of model-free RL, where robust least-square regression is often employed for value function estimation. However, these techniques cannot be directly applied to model-based RL. In this paper, we focus on model-based RL and take the maximum likelihood estimation (MLE) approach to learn transition model. Our work encompasses both online and offline settings. In the online setting, we introduce an algorithm called corruption-robust optimistic MLE (CR-OMLE), which leverages total-variation (TV)-based information ratios as uncertainty weights for MLE. We prove that CR-OMLE achieves a regret of $\tilde{\mathcal{O}}(\sqrt{T}+C)$, where $C$ denotes the cumulative corruption level after $T$ episodes. We also prove a lower bound to show that the additive dependence on $C$ is optimal. We extend our weighting technique to the offline setting, and propose an algorithm named corruption-robust pessimistic MLE (CR-PMLE). Under a uniform coverage condition, CR-PMLE exhibits suboptimality worsened by $\mathcal{O}(C/n)$, nearly matching the lower bound. To the best of our knowledge, this is the first work on corruption-robust model-based RL algorithms with provable guarantees.
@@ 1 Introduction
Reinforcement learning (RL) seeks to find the optimal policy within an unknown environment associated with rewards and transition dynamics. A representative model for RL is the Markov decision process (MDP) (Sutton and Barto,, 2018). While numerous studies assume static rewards and transitions, the environments in real-world scenarios are often non-stationary and vulnerable to adversarial corruptions. For instance, autonomous vehicles frequently fall victim to misled navigation caused by hacked maps and adversarially contaminated traffic signs (Eykholt et al.,, 2018). Similarly, in smart healthcare systems, an adversary with partial knowledge can easily manipulate patient statuses (Newaz et al.,, 2020). Under this situation, standard RL algorithms often fail to find policies robust to such adversarial corruption. Therefore, how to identify the optimal policies against the adversarial corruption has witnessed a flurry of recent investigations.
@@
In this work, we focus on the scenario where the adversary can manipulate the transitions before the agent observes the next state. Achieving a sub-linear regret bound under transition corruption has been shown to be computationally challenging with full information feedback (Abbasi Yadkori et al.,, 2013), and even information-theoretically challenging with bandit feedback (Tian et al.,, 2021). Therefore, a series of studies have introduced constraints on the level of corruption, such as limiting the fraction of corrupted samples (Zhang et al.,, 2022) or the cumulative sum of corruptions over $T$ rounds (Lykouris et al.,, 2018; Gupta et al.,, 2019; He et al.,, 2022; Ye et al., 2023a, ; Ye et al., 2023b, ). While these works exhibit a sub-linear regret bound, to the best of our knowledge, existing works all focus on the setting of model-free RL, where the agent directly learns a policy or a value function from the experiences gained through interactions.
@@
In contrast to model-free RL, in model-based RL, the agent learns an explicit model of the environment and utilizes this model for decision-making. This paradigm has demonstrated impressive sample efficiency in both theory (Kearns and Singh,, 2002; Brafman and Tennenholtz,, 2002; Auer et al.,, 2008; Sun et al.,, 2019; Agarwal and Zhang,, 2022) and applications (Chua et al.,, 2018; Nagabandi et al.,, 2020; Schrittwieser et al.,, 2020). In the online setting, one of the most representative frameworks is Optimistic Maximum Likelihood (OMLE) (Liu et al., 2023a, ). This framework establishes a confidence set based on log-likelihood and selects the most optimistic model within the confidence set. However, how to make model-based RL provably robust against adversarial corruption remains an open problem.
@@
In this paper, we resolve this open problem in both online and offline settings with general function approximation. For simplicity, we assume that the reward is known and mainly focus on learning the transition, which is both unknown and subject to corruption. In particular, we first introduce a cumulative measure for the corruption level of the transition probabilities.
@@
In the online setting, to enhance the resilience of exploitation and exploration to potential corruption, we integrate OMLE with a novel uncertainty weighting technique. Distinct from previous works (He et al.,, 2022; Ye et al., 2023a, ; Ye et al., 2023b, ) in the bandits or model-free RL that characterize uncertainty with rewards or value functions, we characterize the uncertainty with the probability measure of transitions. More specifically, we define the uncertainty as a total-variation (TV)-based information ratio (IR) between the current sample against historical samples. Notably, the introduced IR resembles to the eluder coefficient in Zhang, (2023). The samples with higher uncertainty is down-weighted since they are more vulnerable to corruption. Additionally, similar to Liu et al., 2023a , we quantify the complexity of the model class $\mathcal{M}$ with $TV$-based eluder dimension, and establish a connection between the TV-based eluder dimension and the cumulative TV-based IR following Ye et al., 2023a (by considering uncertainty weights).
@@
  • For the online setting, we propose an algorithm CR-OMLE (Corruption-Robust Optimistic MLE) by integrating uncertainty weights with OMLE. This algorithm enjoys a regret bound of $\tilde{\mathcal{O}}(H\log B\sqrt{T\log|\mathcal{M}|{\mathrm{ED}}}+CH\cdot{\mathrm{ED}})$, where $H$ is the episode length, $B$ is an upper bound of transition ratios, $T$ is the number of episodes, $|\mathcal{M}|$ is the cardinality of the model class, ${\mathrm{ED}}$ is the eluder dimension, and $C$ is the corruption level over $T$ episodes. Moreover, we construct a novel lower bound of $\Omega(HdC)$ for linear MDPs (Jin et al.,, 2020) with dimension $d$. As a result, the corruption-dependent term in the upper bound matches the lower bound, when reduced to the linear setting. These results collectively suggest that our algorithm is not only robust to adversarial corruption but also near-optimal with respect to the corruption level $C$.
  • For the offline setting, we propose an algorithm CR-PMLE (Corruption-Robust Pessimistic MLE) and demonstrate that, given a corrupted dataset with an uniform coverage condition, the suboptimality of the policy generated by this algorithm against the optimal policy can be upper bounded by $\tilde{\mathcal{O}}(H\log B\sqrt{\log|\mathcal{M}|/(T\mathrm{Cov}(\mathcal{M},\mathcal{D}))}+CH/(T\mathrm{Cov}(\mathcal{M},\mathcal{D})))$, where $T$ is the number of trajectories, $\mathrm{Cov}(\mathcal{M},\mathcal{D})$ is the coverage coefficient with respect to model class $\mathcal{M}$ and dataset $\mathcal{D}$. Furthermore, we establish a lower bound of $\Omega\big{(}{C}/{(T\cdot\mathrm{Cov}(\mathcal{M},\mathcal{D})})\big{)}$ for the suboptimality within the linear MDP region. Remarkably, the corruption-dependent term in the upper bound aligns closely with the lower bound.
@@ 2 Related Work
Corruption-Robust Bandits and RL. The adversarial corruption is first brought into multi-armed bandit problems by Lykouris et al., (2018), where rewards face attacks $c_{t}$ in each round, with the cumulative corruption level over $T$ rounds represented as $C=\sum_{t=1}^{T}|c_{t}|$. A regret lower bound of $o(T)+O(C)$ was constructed by Gupta et al., (2019), illustrating that an “entangled” relationship between $T$ and $C$ is ideal. Subsequently, the corruption was extended to linear contextual bandits by Bogunovic et al., (2021); Ding et al., (2022); Foster et al., (2020); Lee et al., (2021); Zhao et al., (2021); Kang et al., (2023). However, these approaches either obtains sub-optimal regret bounds or necessitate additional assumptions. He et al., (2022) first achieved a regret bound that matches the lower bound by utilizing a uncertainty weighting technique. Turning to MDPs, the corruption in transitions triggers considerable interest. Wu et al., (2021) first studied corruption on rewards and corruption simultaneously for tabular MPDs. Later, a unified framework to deal with unknown corruption was established by Wei et al., (2022), involving a weak adversary who decides the corruption amount for each action before observing the agent’s action. Subsequently, Ye et al., 2023a ; Ye et al., 2023b extended the uncertainty weighting technique from He et al., (2022) to RL with general function approximation for both and offline settings respectively, aligning with the lower bound in terms of the corruption-related term. However, previous works all focused on the model-free setting, where models are learned via regression. Notably, corruption in the model-based setting, where the models are learned via maximum likelihood estimation (MLE), remains an unexplored area in the literature.
@@
Model-Based RL. There is an emerging body of literature addressing model-based RL problems, from tabular MDPs (Kearns and Singh,, 2002; Brafman and Tennenholtz,, 2002; Auer et al.,, 2008) to rich observation spaces with linear function approximation (Yang and Wang,, 2020; Ayoub et al.,, 2020). Extending beyond linear settings, recent studies (Sun et al.,, 2019; Agarwal et al.,, 2020; Uehara et al.,, 2021; Du et al.,, 2021; Agarwal and Zhang,, 2022; Wang et al.,, 2023; Liu et al., 2023a, ; Foster et al.,, 2021; Zhong et al., 2022b, ; Chen et al.,, 2022) have explored general function approximations. Simultaneously, a parallel line of research has focused on the model-free approach with general function approximations (Jin et al.,, 2021; Jiang et al.,, 2017; Chen et al.,, 2022; Zhang,, 2023; Agarwal et al.,, 2023; Zhao et al.,, 2023; Di et al.,, 2023; Liu et al., 2023b, ). Among the works, the most relevant one to ours is Liu et al., 2023a , who makes optimistic estimations via a log-likelihood approach and generalizes an eluder-type condition from Russo and Van Roy, (2013). Despite numerous works on online model-based RL problems, to the best of our knowledge, there is still a gap in the literature regarding a theoretical guarantee for offline model-based RL.
@@ 3 Preliminaries
In model-based RL, we consider an MDP model class $\mathcal{M}$. Without loss of generality, we assume that the class $\mathcal{M}$ has finite elements. Each model $M\in\mathcal{M}$ depicts transition probability $P_{M}^{h}(x^{h+1}|x,a)$. We use ${\mathbb{E}}^{M}[\cdot]$ to represent the expectation over the trajectory under transition probability $P_{M}$, $V_{M}$ and $Q_{M}$ to denote the $V$-value and $Q$-value function of model $M$, and $\pi_{M}=\pi_{f_{M}}$ to denote the greedy policy for value function $f_{M}$. Additionally, we suppose that there exists an underlying true model $M^{*}\in\mathcal{M}$ such that the true transition probability is $P_{M_{*}}^{h}(x^{h+1}|x^{h},a^{h})$, and use the short-hand notation $P_{*}^{h}=P_{M^{*}}^{h}$ and ${\mathbb{E}}^{*}[\cdot]={\mathbb{E}}^{M_{*}}[\cdot]$. Given a model $M$, the model-based Bellman error is defined as
$\mathcal{E}^{h}(M,x^{h},a^{h})$ $=Q_{M}^{h}(x^{h},a^{h})-R_{*}^{h}(x^{h},a^{h})-{\mathbb{E}}^{*}[V_{M}^{h+1}(x^{h+1}|x^{h},a^{h}]$
@@
$\sup_{M\in\mathcal{M}}\Big{\|}\frac{P_{M}^{h}}{P_{*}^{h}}\Big{\|}\leq B.$
@@
To depict the structure of the model space $\mathcal{M}$, we introduce the model-based version of the eluder dimension similar to Russo and Van Roy, (2013); Liu et al., 2023a .
@@
For an $\epsilon>0$, we say that a point $z$ is $\epsilon$-dependent on a set $\mathcal{Z}$ with respect to $\mathcal{M}$ if there exists $M,M^{\prime}\in\mathcal{M}$ such that $\sum_{s=1}^{i-1}|{\mathrm{TV}}(P^{h}_{M_{i}}(\cdot|z_{s})\|P^{h}_{M_{i}^{\prime}}(\cdot|z_{s}))|^{2}\leq\epsilon^{2}$ implies $|{\mathrm{TV}}(P^{h}_{M}(\cdot|z)\|P^{h}_{M_{i}^{\prime}}(\cdot|z_{i}))|\leq\epsilon$.
@@
For a model class $\mathcal{M}$, an $\epsilon>0$, the ${\mathrm{TV}}$-eluder dimension ${\mathrm{ED}}(\mathcal{M},\epsilon)$ is the length $n$ of the longest sequence $\{z_{1},\ldots,z_{n}\}\subset\mathcal{X}\times\mathcal{A}$ such that for some $\epsilon^{\prime}\geq\epsilon$ and all $h\in[H]$, $i\in[n]$, $z_{i}$ is ϵsuperscriptitalic-$\epsilon^{\prime}$ independent of its predecessors.
Particularly, Liu et al., 2023a also defines the TV-eluder condition. The distinction lies in their consideration of a sequence of policies, while we focus on a sequence of state-action samples. In Theorems D.1 and D.2, we offer examples of tabular and linear MDPs where the TV-eluder dimension can be bounded. To facilitate analysis, we formulate the TV-norm version of the eluder coefficient drawing inspiration from the approach of Zhang, (2023). Given a model class $\mathcal{M}^{\prime}$ and sample set $\mathcal{S}_{t}^{h}=\{z_{s}^{h}\}_{s=1}^{t}$, the information ratio (IR) between the estimation error and the training error within $\mathcal{M}^{\prime}$ with respect to $\bar{M}_{t}$ is defined as
@@
where $l_{t}^{h}(M,M^{\prime})$ denotes ${\mathrm{TV}}(P_{M^{\prime}}^{h}(\cdot|z_{t}^{h})\|P_{M^{\prime}}^{h}(\cdot|z_{t}^{h}))$. We use the linear MDP as an example to illustrate IR. If the transition can be embedded as $P^{h}_{M}(x^{h+1}|z^{h})=\nu^{h}(M,x^{h+1})^{\top}\phi^{h}(z^{h})$, according to Example D.1, the IR can be reduced to
$\min\Big{\{}1,\|\phi^{h}(z_{t}^{h})\|_{(\Sigma_{t}^{h})^{-1}}\Big{\}},$
where $\Sigma_{t}^{h}=\sum_{s=1}^{t-1}\phi^{h}(z_{s}^{h})\phi^{h}(z_{s}^{h})^{\top}$. Intuitively, this quantity represents the uncertainty of vector $\nu^{h}(M,x^{h+1})$ in the direction of $\phi(z_{t}^{h})$. After observing the state-action pair $(x_{t}^{h},a_{t}^{h})$ in each round, an adversary corrupts the transition dynamics from $P_{*}^{h}$ to $P_{t}^{h}$ and the agent receive the next state $x_{t}^{h+1}$ with according to $P_{t}^{h}(\cdot|x_{t}^{h},a_{t}^{h})$. To measure the corruption level, we define the cumulative corruption as follows.

### v3

@@ Abstract
This study tackles the challenges of adversarial corruption in model-based reinforcement learning (RL), where the transition dynamics can be corrupted by an adversary. Existing studies on corruption-robust RL mostly focus on the setting of model-free RL, where robust least-square regression is often employed for value function estimation. However, the uncertainty weighting techniques cannot be directly applied to model-based RL. In this paper, we focus on model-based RL and take the maximum likelihood estimation (MLE) approach to learn transition model. Our work encompasses both online and offline settings. In the online setting, we introduce an algorithm called corruption-robust optimistic MLE (CR-OMLE), which leverages total-variation (TV)-based information ratios as uncertainty weights for MLE. We prove that CR-OMLE achieves a regret of $\tilde{\mathcal{O}}(\sqrt{T}+C)$, where $C$ denotes the cumulative corruption level after $T$ episodes. We also prove a lower bound to show that the additive dependence on $C$ is optimal. We extend our weighting technique to the offline setting, and propose an algorithm named corruption-robust pessimistic MLE (CR-PMLE). Under a uniform coverage condition, CR-PMLE exhibits suboptimality worsened by $\mathcal{O}(C/n)$, nearly matching the lower bound. To the best of our knowledge, this is the first work on corruption-robust model-based RL algorithms with provable guarantees.
Machine Learning, ICML
@@ 1 Introduction
Reinforcement learning (RL) seeks to find the optimal policy within an unknown environment associated with rewards and transition dynamics. A representative model for RL is the Markov decision process (MDP) (Sutton & Barto, 2018). While numerous studies assume static rewards and transitions, the environments in real-world scenarios are often non-stationary and vulnerable to adversarial corruption. For instance, autonomous vehicles frequently fall victim to misled navigation caused by hacked maps and adversarially contaminated traffic signs (Eykholt et al., 2018). Similarly, in smart healthcare systems, an adversary with partial knowledge can easily manipulate patient statuses (Nanayakkara et al., 2022). Under this situation, standard RL algorithms often fail to find policies robust to such adversarial corruption. Therefore, how to identify the optimal policies against adversarial corruption has witnessed a flurry of recent investigations.
@@
In this work, we focus on the scenario where the adversary can manipulate the transitions before the agent observes the next state. Achieving a sub-linear regret bound under transition corruption has been shown to be computationally challenging with full information feedback (Abbasi Yadkori et al., 2013), and even information-theoretically challenging with bandit feedback (Tian et al., 2021). Therefore, a series of studies have introduced constraints on the level of corruption, such as limiting the fraction of corrupted samples (Zhang et al., 2022) or the cumulative sum of corruptions over $T$ rounds (Lykouris et al., 2018; Gupta et al., 2019; He et al., 2022; Ye et al., 2023a, b; Yang et al., 2023). While these works exhibit a sub-linear regret bound, to the best of our knowledge, existing works all focus on the setting of model-free RL, where the agent directly learns a policy or a value function from the experiences gained through interactions.
@@
In contrast to model-free RL, in model-based RL, the agent learns an explicit model of the environment and utilizes this model for decision-making. This paradigm not only exempts from Bellman completeness (Jin et al., 2021) (which is a standard assumption for model-free RL) but also has demonstrated impressive sample efficiency in both theories (Kearns & Singh, 2002; Brafman & Tennenholtz, 2002; Auer et al., 2008; Sun et al., 2019; Agarwal & Zhang, 2022) and applications (Chua et al., 2018; Nagabandi et al., 2020; Schrittwieser et al., 2020), such as robotics (Polydoros & Nalpantidis, 2017) and autonomous driving (Wu et al., 2021). In the online setting, one of the most representative frameworks is Optimistic Maximum Likelihood (OMLE) (Liu et al., 2023a). This framework establishes a confidence set based on log-likelihood and selects the most optimistic model within the confidence set. However, how to make model-based RL provably robust against adversarial corruption remains an open problem.
@@
In this paper, we resolve this open problem in both online and offline settings with a general function approximation. For simplicity, we assume that the reward is known and mainly focuses on learning the transition, which is both unknown and subject to corruption. In particular, we first introduce a cumulative measure for the corruption level of the transition probabilities.
@@
In the online setting, to enhance the resilience of exploitation and exploration to potential corruption, we integrate OMLE with a novel uncertainty weighting technique. Distinct from previous works (He et al., 2022; Ye et al., 2023a, b) in the bandits or model-free RL that characterize uncertainty with rewards or value functions, we characterize the uncertainty with the probability measure of transitions. More specifically, we define the uncertainty as a total-variation (TV)-based information ratio (IR) between the current sample and historical samples. Notably, the introduced IR resembles the eluder coefficient in Zhang (2023). The samples with higher uncertainty are down-weighted since they are more vulnerable to corruption. Additionally, similar to Liu et al. (2023a), we quantify the complexity of the model class $\mathcal{M}$ with $TV$-based eluder dimension and establish a connection between the TV-based eluder dimension and the cumulative TV-based IR following Ye et al. (2023a) (by considering uncertainty weights).
@@
  • For the online setting, we propose an algorithm CR-OMLE (Corruption-Robust Optimistic MLE) by integrating uncertainty weights with OMLE. This algorithm enjoys a regret bound of $\tilde{\mathcal{O}}(H\log B\sqrt{T\log|\mathcal{M}|{\mathrm{ED}}}+CH\cdot{\mathrm{ED}})$, where $H$ is the episode length, $B$ is an upper bound of transition ratios, $T$ is the number of episodes, $|\mathcal{M}|$ is the cardinality of the model class, ${\mathrm{ED}}$ is the eluder dimension, and $C$ is the corruption level over $T$ episodes. Moreover, we construct a novel lower bound of $\Omega(HdC)$ for linear MDPs (Jin et al., 2020) with dimension $d$. As a result, the corruption-dependent term in the upper bound matches the lower bound, when reduced to the linear setting. These results collectively suggest that our algorithm is not only robust to adversarial corruption but also near-optimal with respect to the corruption level $C$.
  • For the offline setting, we propose an algorithm CR-PMLE (Corruption-Robust Pessimistic MLE) and demonstrate that, given a corrupted dataset with an uniform coverage condition, the suboptimality of the policy generated by this algorithm against the optimal policy can be upper bounded by $\tilde{\mathcal{O}}(H\log B\sqrt{\log|\mathcal{M}|/(T\mathrm{Cov}(\mathcal{M},\mathcal{D}))}+CH/(T\mathrm{Cov}(\mathcal{M},\mathcal{D})))$, where $T$ is the number of trajectories, $\mathrm{Cov}(\mathcal{M},\mathcal{D})$ is the coverage coefficient with respect to model class $\mathcal{M}$ and dataset $\mathcal{D}$. Furthermore, we establish a lower bound of $\Omega\big{(}{C}/{(T\cdot\mathrm{Cov}(\mathcal{M},\mathcal{D})})\big{)}$ for the suboptimality within the linear MDP region. Remarkably, the corruption-dependent term in the upper bound aligns closely with the lower bound up to $H$. It is also worth noting that this is also the first provable model-based offline RL algorithm even without considering the corruption.
@@ 2 Related Work
Corruption-Robust Bandits and RL. The adversarial corruption is first brought into multi-armed bandit problems by Lykouris et al. (2018), where rewards face attacks by $c_{t}$ in each round, with the cumulative corruption level over $T$ rounds represented as $C=\sum_{t=1}^{T}|c_{t}|$. A regret lower bound of $o(T)+O(C)$ was constructed by Gupta et al. (2019), illustrating that an “entangled” relationship between $T$ and $C$ is ideal. Subsequently, the corruption was extended to linear contextual bandits by Bogunovic et al. (2021); Ding et al. (2022); Foster et al. (2020); Lee et al. (2021); Zhao et al. (2021); Kang et al. (2023). However, these approaches either obtain sub-optimal regret bounds or necessitate additional assumptions. He et al. (2022) first achieved a regret bound that matches the lower bound by utilizing an uncertainty weighting technique. Turning to MDPs, the corruption in transitions triggers considerable interest. Wu et al. (2021) first studied corruption on rewards and corruption simultaneously for tabular MPDs. Later, a unified framework to deal with unknown corruption was established by Wei et al. (2022), involving a weak adversary who decides the corruption amount for each action before observing the agent’s action. Subsequently, Ye et al. (2023a, b) extended the uncertainty weighting technique from He et al. (2022) to RL with general function approximation for both online and offline settings respectively, aligning with the lower bound in terms of the corruption-related term. However, previous works all focused on the model-free setting, where models are learned via regression. Notably, corruption in the model-based setting, where the models are learned via maximum likelihood estimation (MLE), remains an unexplored area in the literature.
@@
Model-Based RL. There is an emerging body of literature addressing model-based RL problems, from tabular MDPs (Kearns & Singh, 2002; Brafman & Tennenholtz, 2002; Auer et al., 2008) to rich observation spaces with linear function approximation (Yang & Wang, 2020; Ayoub et al., 2020). Extending beyond linear settings, recent studies (Sun et al., 2019; Agarwal et al., 2020; Uehara et al., 2021; Du et al., 2021; Agarwal & Zhang, 2022; Wang et al., 2023; Liu et al., 2023a; Foster et al., 2021; Zhong et al., 2022b; Chen et al., 2022) have explored general function approximations. Simultaneously, a parallel line of research has focused on the model-free approach with general function approximations (Jin et al., 2021; Jiang et al., 2017; Chen et al., 2022; Zhang, 2023; Agarwal et al., 2023; Zhao et al., 2023; Di et al., 2023; Liu et al., 2023b). Among the works, the most relevant one to ours is Liu et al. (2023a), who makes optimistic estimations via a log-likelihood approach and generalizes an eluder-type condition from Russo & Van Roy (2013). Despite numerous works on online model-based RL problems, to the best of our knowledge, there is still a gap in the literature regarding a theoretical guarantee for offline model-based RL.
Distributional Robust RL. Our work shares the same goal of robustness with another line of work, called distributional robust RL, which formulates the uncertainty of the transitions as an uncertainty set (Roy et al., 2017; Badrinath & Kalathil, 2021; Wang & Zou, 2021) or a set of distributions surrounding the nominal models (Zhou et al., 2021; Shi et al., 2024; Clavier et al., 2023). In this setting, there is no corruption during the training process or exploration phase, and the aim is to find a robust policy that maintains a near-optimal policy when there exists a small distributional shift between the training and real environments. In comparison, our work focuses on cases where the training data is corrupted, and we aim to identify the optimal policy for the hidden environment even with a few corrupted observations. Hence, these two categories of works study different notions of robustness and have different challenges.
@@ 3 Preliminaries
In model-based RL, we consider an MDP model class $\mathcal{M}$. Without loss of generality, we assume that the class $\mathcal{M}$ has finite elements. Note that the finite class assumption is only to simplify the analysis. We can extend the proof to an infinite model class by taking a finite covering. Each model $M\in\mathcal{M}$ depicts transition probability $P_{M}^{h}(x^{h+1}|x,a)$. We use ${\mathbb{E}}^{M}[\cdot]$ to represent the expectation over the trajectory under transition probability $P_{M}^{\pi}$, $V_{M}^{\pi}$ and $Q_{M}^{\pi}$ to denote the $V$-value and $Q$-value function of model $M$ and policy $\pi$, and $\pi_{M}=\pi_{Q_{M}}$ to denote the optimal policy for the model $M$. Then, we short-notate $V_{M}^{\pi_{M},h}$ and $Q_{M}^{\pi_{M},h}$ as $V_{M}^{h}$ and $Q_{M}^{h}$. Additionally, we suppose that there exists an underlying true model $M^{*}\in\mathcal{M}$ such that the true transition probability is $P_{M_{*}}^{h}(x^{h+1}|x^{h},a^{h})$, and use the short-hand notation $P_{*}^{h}=P_{M^{*}}^{h}$ and ${\mathbb{E}}^{*}[\cdot]={\mathbb{E}}^{M_{*}}[\cdot]$. Given a model $M$, the model-based Bellman error is defined as
$\mathcal{E}^{h}(M,x^{h},a^{h})$
$=Q_{M}^{h}(x^{h},a^{h})-R_{*}^{h}(x^{h},a^{h})-{\mathbb{E}}^{*}[V_{M}^{h+1}(x^{h+1})|x^{h},a^{h}]$
@@
$\sup_{h\in[H],M\in\mathcal{M}}\max\bigg{\{}\Big{\|}\frac{P_{*}^{h}}{P_{M}^{h}}\Big{\|}_{\infty}\,\Big{\|}\frac{P_{M}^{h}}{P_{*}^{h}}\Big{\|}_{\infty}\bigg{\}}\leq B.$
@@
To depict the structure of the model space $\mathcal{M}$, we introduce the model-based version of the eluder dimension similar to Russo & Van Roy (2013); Liu et al. (2023a).
@@
For an $\epsilon>0$, we say that a point $z$ is $\epsilon$-dependent on a set $\mathcal{Z}$ with respect to $\mathcal{M}$ if there exists $M,M^{\prime}\in\mathcal{M}$ such that $\sum_{z^{\prime}\in\mathcal{Z}}|\mathrm{tv}(P^{h}_{M}(\cdot|z^{\prime})\|P^{h}_{M^{\prime}}(\cdot|z^{\prime}))|^{2}\leq\epsilon^{2}$ implies $|\mathrm{tv}(P^{h}_{M}(\cdot|z)\|P^{h}_{M^{\prime}}(\cdot|z))|\leq\epsilon$.
@@
For a model class $\mathcal{M}$, an $\epsilon>0$, the ${\mathrm{TV}}$-eluder dimension ${\mathrm{ED}}(\mathcal{M},\epsilon)$ is the length $n$ of the longest sequence $\{z_{1},\ldots,z_{n}\}\subset\mathcal{X}\times\mathcal{A}$ such that for some $\epsilon^{\prime}\geq\epsilon$ and all $h\in[H]$, $i\in[n]$, $z_{i}$ is $\epsilon^{\prime}$ independent of its predecessors.
Particularly, Liu et al. (2023a) also defines the TV-eluder condition. The distinction lies in their consideration of a sequence of policies, while we focus on a sequence of state-action samples. In Theorems D.1 and D.2, we offer examples of tabular and linear MDPs where the TV-eluder dimension can be bounded. To facilitate analysis, we formulate the TV-norm version of the eluder coefficient drawing inspiration from the approach of Zhang (2023). Given a model class $\mathcal{M}^{\prime}$, sample set $\mathcal{S}_{t}^{h}=\{z_{s}^{h}\}_{s=1}^{t}$ and some estimator $\bar{M}_{t}$ from $\mathcal{S}_{t}^{h}$ (will be specified by later algorithms), the information ratio (IR) between the estimation error and the training error within $\mathcal{M}^{\prime}$ with respect to $\bar{M}_{t}$ is
@@
where $l_{t}^{h}(M,M^{\prime})$ denotes ${\mathrm{TV}}(P_{M}^{h}(\cdot|z_{t}^{h})\|P_{M^{\prime}}^{h}(\cdot|z_{t}^{h}))$. We use the linear MDP as an example to illustrate IR. If the transition can be embedded as $P^{h}_{M}(x^{h+1}|z^{h})=\nu^{h}(M,x^{h+1})^{\top}\phi^{h}(z^{h})$, according to Example D.3, the IR can be reduced to
$\min\Big{\{}1,\|\phi^{h}(z^{h})\|_{(\Sigma_{t}^{h})^{-1}}\Big{\}},$
where $\Sigma_{t}^{h}=\sum_{s=1}^{t-1}\phi^{h}(z_{s}^{h})\phi^{h}(z_{s}^{h})^{\top}$. Intuitively, this quantity represents the uncertainty of vector $\nu^{h}(M,x^{h+1})$ in the direction of $\phi(z_{t}^{h})$. After observing the state-action pair $(x_{t}^{h},a_{t}^{h})$ in each round, an adversary corrupts the transition dynamics from $P_{*}^{h}$ to $P_{t}^{h}$ and the agent receives the next state $x_{t}^{h+1}$ induced by $P_{t}^{h}(\cdot|x_{t}^{h},a_{t}^{h})$. To measure the corruption level, we define cumulative corruption.